You are a project manager for a manufacturing company. One of the machine parts on the assembly line is a thin, uniform rod that is60.0 cm long and has mass 0.900 kg .

Part A:What is the moment of inertia of this rod for an axis at its center, perpendicular to the rod?

Part B: One of your engineers has proposed to reduce the moment of inertia by bending the rod at its center into a V-shape, with a 60.0 degree angle at its vertex. What would be the moment of inertia of this bent rod about an axis perpendicular to the plane of the V at its vertex?
Answer #1

Part A) I = ML2/12 = 0.9 * 0.62 / 12 = 0.027 kg-m2

Part B) I = 2[(M/2)(L/2)2/3] = ML2/12 = 0.9 * 0.62 / 12 = 0.027 kg-m2